# Contrastive Multi-Task Learning with Solvent-Aware Augmentation for Drug Discovery

### v1

@@ III-A2 Solvent-Aware Conformer Generation
Methanol 1 107 5121\,107\,5121 107 512 Hexafluoroacetone 367 490367\,490367 490
DMSO 1 105 4471\,105\,4471 105 447 HMPA 367 365367\,365367 365
Chloroform 1 089 6961\,089\,6961 089 696 DCM 367 312367\,312367 312
Butylformate 370 327370\,327370 327 Benzonitrile 367 008367\,008367 008
NMP 370 315370\,315370 315 Hexafluorobenzene 366 831366\,831366 831
Nitromethane 370 304370\,304370 304 Propionitrile 366 828366\,828366 828
Sulfolane 370 280370\,280370 280 Pyridine 366 067366\,067366 067
IPA 370 158370\,158370 158 Carbon tetrachloride 365 678365\,678365 678
DMPU 369 745369\,745369 745 Acetonitrile 365 464365\,464365 464
Octanol 369 453369\,453369 453 Cyclohexane 365 364365\,364365 364
Ethyl acetate 369 272369\,272369 272 Benzene 364 541364\,541364 541
Glycerin 368 570368\,570368 570 1,4-Dioxane 364 334364\,334364 334
DMF 367 528367\,528367 528 Acetone 364 183364\,183364 183
Diethyl ether 364 052364\,052364 052 2-Nitropropane 359 196359\,196359 196
Nitrobenzene 363 863363\,863363 863 MTBE 359 136359\,136359 136
Trifluorotoluene 363 727363\,727363 727 Acetic acid 359 122359\,122359 122
THF 361 535361\,535361 535 Oxylol 358 606358\,606358 606
Toluene 361 517361\,517361 517 DME 308 870308\,870308 870
Hexane 361 073361\,073361 073 Ethanol 360 079360\,079360 079
@@ IV-A Performance Comparison in Downstream Tasks
IV-A1 Downstream Task 1: Ligand Binding Affinity
@@
IV-A2 Downstream Task 2: Molecular Docking
@@
Our model achieves the highest success rate among all evaluated methods when the RMSD threshold is set at $<$ 5.0 Å in CASF-2016 (Table III). In the more stringent PoseBusters (428 complexes) and Astex (85 complexes) benchmarks [36], it delivers 64% and 82% success rates, respectively, at the $<$ 2.0 Å cutoff, shown in Fig. 4. Notably, since Uni-Mol Docking v2 was trained on a large amount of data with a training time three times longer than ours, these results demonstrate that our model can achieve competitive performance even after fine-tuning in a short period.
@@
Figure 4: Performance comparison on PoseBusters and Astex. (a) overall success rates ($<$2.0 Å RMSD) across 428 PoseBusters complexes and 85 Astex entries for all evaluated methods; “Ours” is highlighted. (b) case-study of structural superposition for PDB entry 2J7H (Thermotoga maritima $\beta$-glucosidase bound to azafagomine), showing the predicted pose (green) overlaid on the crystallographic pose (red).
IV-A3 Downstream Task 3: Virtual Screening
@@ IV-B Ablation Study
The synergistic integration of objectives for geometric self-supervised learning establishes state-of-the-art performance. MMR provides the essential foundation through solvent-contextualized local geometric reconstruction, explicitly modeling hydration effects at binding interfaces and enabling precise torsion angle prediction via continuous geometric learning. Its removal induces significant performance variance (44.1%), underscoring its necessity. IDMP delivers complementary global geometric constraints by learning pairwise atomic distances across ligand states, which enhances prediction robustness (AUC) while reducing variance. Contrastive learning emerges as the critical performance amplifier and stabilizer. When integrated with MMR+IDMP, CL reduces cross-validation variance by 46.8%.
The integrated framework demonstrates superior reliability, with AUC variance reduced by 18.3% and ROC enrichment at 5% variance by 25.3% versus MMR-only. CL emerges as the principal stabilizer through noise-invariant representation learning, while GCE loss enhances geometric reliability.
@@ V Conclusion
ACKNOWLEDGEMENTS
This work was supported by an internal grant from The Hong Kong Polytechnic University (Project No. P0051278, Jung Sun Yoo) and the General Research Fund (Project No. PolyU 15101422, Jung Sun Yoo) from the Research Grants Council of the Hong Kong Special Administrative Region, China.
@@ V-A8 $\boldsymbol{r_{m}^{2}}$ Coefficient of Determination
V-B Downstream Task: Drug-Target Affinity Prediction
To demonstrate versatility, we extend evaluation to the Drug–Target Affinity (DTA) task using TDC’s DAVIS (Kd values; 25,772 interactions) and KIBA (KIBA scores; 117,657 interactions) datasets. We also utilize auxiliary binding site data ( 18M protein sequence and ChEBI ID pairs). For DTA, we employ the Concordance Index (CI) and the modified coefficient of determination ($r_{m}^{2}$), considering $r_{m}^{2}\geq 0.5$ acceptable, as shown in Table VI.
TABLE VI: Performance comparison on DAVIS and KIBA benchmarks. Best in bold; second-best underlined.
Method DAVIS KIBA
MSE↓ CI↑ $r_{m}^{2}$ MSE↓ CI↑ $r_{m}^{2}$
KronRLS 0.379 0.871 0.407 0.411 0.782 0.342
SimBoost 0.282 0.872 0.644 0.222 0.836 0.629
SimCNN-DTA 0.319 0.852 0.595 0.274 0.821 0.573
DeepDTA 0.261 0.878 0.630 0.194 0.863 0.673
WideDTA 0.886 0.262 0.875 0.179
AttentionDTA 0.216 0.893 0.677 0.155 0.882 0.755
MATT-DTI 0.227 0.891 0.683 0.150 0.882 0.756
GraphDTA 0.229 0.893 0.649 0.139 0.889 0.736
FusionDTA 0.220 0.903 0.666 0.167 0.891 0.699
BiCompDTA 0.237 0.904 0.696 0.167 0.891 0.757
GANsDTA 0.276 0.881 0.653 0.224 0.866 0.775
TF-DTA 0.231 0.886 0.670 0.177 0.877 0.734
MGraphDTA 0.207 0.900 0.710 0.128 0.902 0.801
DGraphDTA 0.202 0.904 0.700 0.126 0.904 0.786
WGNN-DTA 0.208 0.900 0.692 0.144 0.885 0.781
PocketDTA 0.177 0.903 0.731 0.140 0.892 0.771
MMD-DTA 0.220 0.905 0.134 0.900
AffinityLM 0.205 0.899 0.712 0.140 0.899 0.789
Ours 0.238 0.883 0.665 0.144 0.902 0.789

### v2

@@ III-A2 Solvent-Aware Conformer Generation
Methanol 1 107 5121\,107\,512 Hexafluoroacetone 367 490367\,490
DMSO 1 105 4471\,105\,447 HMPA 367 365367\,365
Chloroform 1 089 6961\,089\,696 DCM 367 312367\,312
Butylformate 370 327370\,327 Benzonitrile 367 008367\,008
NMP 370 315370\,315 Hexafluorobenzene 366 831366\,831
Nitromethane 370 304370\,304 Propionitrile 366 828366\,828
Sulfolane 370 280370\,280 Pyridine 366 067366\,067
IPA 370 158370\,158 Carbon tetrachloride 365 678365\,678
DMPU 369 745369\,745 Acetonitrile 365 464365\,464
Octanol 369 453369\,453 Cyclohexane 365 364365\,364
Ethyl acetate 369 272369\,272 Benzene 364 541364\,541
Glycerin 368 570368\,570 1,4-Dioxane 364 334364\,334
DMF 367 528367\,528 Acetone 364 183364\,183
Diethyl ether 364 052364\,052 2-Nitropropane 359 196359\,196
Nitrobenzene 363 863363\,863 MTBE 359 136359\,136
Trifluorotoluene 363 727363\,727 Acetic acid 359 122359\,122
THF 361 535361\,535 Oxylol 358 606358\,606
Toluene 361 517361\,517 DME 308 870308\,870
Hexane 361 073361\,073 Ethanol 360 079360\,079
@@ IV-A Performance Comparison in Downstream Tasks
IV-A1 Downstream Task 1. Ligand Binding Affinity
@@
IV-A2 Downstream Task 2. Molecular Docking
@@
Our model achieves the highest success rate among all evaluated methods when the RMSD threshold is set at << 5.0 Å in CASF-2016 (Table III). In the more stringent PoseBusters (428 complexes) and Astex (85 complexes) benchmarks [36], it delivers 64% and 82% success rates, respectively, at the << 2.0 Å cutoff, shown in Fig. 4. Notably, since Uni-Mol Docking v2 was trained on a large amount of data with a training time three times longer than ours, these results demonstrate that our model can achieve competitive performance even after fine-tuning in a short period.
@@
Figure 4: Performance comparison on PoseBusters and Astex. (a) overall success rates (<<2.0 Å RMSD) across 428 PoseBusters complexes and 85 Astex entries for all evaluated methods; “Ours” is highlighted. (b) case-study of structural superposition for PDB entry 2J7H (Thermotoga maritima $\beta$-glucosidase bound to azafagomine), showing the predicted pose (green) overlaid on the crystallographic pose (red).
IV-A3 Downstream Task 3. Virtual Screening
@@ IV-B Ablation Study
The results in Table V demonstrate the individual and joint contributions of each architectural module. Notably, the full model, incorporating MMR, IDMP, and CL, achieves the highest performance across all metrics, with an AUC of $97.14\pm 0.0098\%$ and substantial gains in early enrichment. This highlights the significance of combining local, global, and solvent-invariant representations.
Contribution of Individual Modules
Ablating the MMR module caused the largest drop in both AUC and enrichment metrics, underscoring its importance for local geometric features. When IDMP was absent yet MMR was maintained, the performance exhibited a slight decline, indicating that global spatial modeling enhances the understanding of molecular interactions. Exclusion of CL consistently reduced early enrichment, especially at 0.5%, confirming its critical role in solvent-invariant robustness.
Synergistic Effects and Model Robustness
Ablation of individual modules consistently lowers performance, confirming the synergistic interaction among the architectural components. Incorporating both IDMP and CL alongside MMR increases the AUC and further reduces performance variance, demonstrating that the integrated design provides enhanced statistical stability and robustness. The consistent improvements observed across all enrichment factor thresholds further attest to the robustness and practical utility of the integrated model in virtual screening applications.